A sample random sample of size 4 is drown with replacement from a population with s.d. 6 what is the S. E of the sample mean?

1
Expert's answer
2022-05-18T12:14:52-0400
"SE=\dfrac{\sigma}{\sqrt{n}}=\dfrac{6}{\sqrt{4}}=3"